OpenStudy (anonymous):

Tyrece's quarter grade in Science is calculated by finding his mean test score for that quarter. So far this quarter, Tyrece has taken three tests and scored 68, 83, and 74. He wants to earn an 80 for the quarter and has his last test on Friday. What grade does Tyrece need to earn on his last test so that he will have an 80 for the quarter?

OpenStudy (anonymous):

80*4=320 320-68-83-74=95 Check: 95+74+83+68=320/4=80